# Detection of the papermilling behavior

### v2

@@ 10 Conclusion
It is obvious that the evaluation of the quality of research performance cannot be based on a single number; it does not matter if that number is the h-index, or the number of publications, or the number of citations, or the integrity index proposed in this paper, or any other single quantitative indicator.
However, there are some standard patterns typical for conscientious researchers, focused on the advancement of knowledge, and for researchers exhibiting the papermilling behavior, focusing on the numbers of publications and citations which give them advantages and benefits.
@@
The detailed analysis of the dynamics of the changes of the number of co-authors related to papermilling behavior requires the change of the format of the data exported from the Web of Science database (or eventually from the Scopus database, which also can be used for such studies).
Acknowledgments
The author is supported by grants APVV-22-0508, VEGA 1/0674/23, KEGA 006TUKE-4/2024, and ARO W911NF-22-1-0264.

### v4

@@ 13 Conclusion
Obviously, the evaluation of the quality of research performance cannot be based on a single number, and it does not matter if that number is the h-index, or the number of publications, or the number of citations, or the integrity index proposed in this paper, or any other single quantitative indicator.
However, there are some standard patterns typical for conscientious researchers, focused on the advancement of knowledge, and for researchers exhibiting the papermilling behavior, focusing on the numbers of publications, number of citations, and h-index, which give them advantages and benefits.
@@
From the presented study it follows that, in addition to the recommendations listed in [3], the journal editors should pay attention not only to each submitted manuscript, avoiding the products of the paper mills, but also to the publication patterns of all its authors.
The detailed analysis of the dynamics of the changes of the number of co-authors related to the papermilling behavior requires the change of the format of the data exported from the Web of Science database (or eventually from the Scopus database, which also can be used for such studies).